228360:

Jeremy finds a set of plans to build a model airplane. Most, but not all, of the lengths given in the plans are direct measurements. According to the plans, each horizontal stabilizer of the tail wing is made from a right triangle. The length of the leg of the triangle that touches the plane is AB = 8 inches, cos=8/14 and . To find the number of inches each horizontal stabilizer extends from the plane, find BC.

Vocaloid:

cosA = adjacent/hypotenuse = 8/14. the adjacent side AB is 8 as stated by the problem, so the hypotenuse must be 14. now we have one leg of a right triangle as well as the hypotenuse, so we can solve for the other leg BC. using the pythagorean theorem (BC)^2 + 8^2 = 14^2 solve for BC